### 03 Jan 2024 v1

# Text mining arXiv: a look through

Abstract. This paper explores articles hosted on the arXiv preprint server with the aim to uncover valuable insights hidden in this vast collection of research. Employing text mining techniques and through the application of natural language processing methods, we examine the contents of quantitative finance papers posted in arXiv from 1997 to 2022. We extract and analyze crucial information from the entire documents, including the references, to understand the topics trends over time and to find out the most cited researchers and journals on this domain. Additionally, we compare numerous algorithms to perform topic modeling, including state-of-the-art approaches.
Key words: quantitative finance, text mining, natural language processing, unsupervised clustering, topic modeling, research trends, named entity recognition.
1 Introduction
Quantitative finance is a field of finance that studies mathematical and statistical models and applies them to financial markets and investments, for pricing, risk management, and portfolio allocation. These models are needed to analyze financial data, to find the price of financial instruments and to measure their risk (see Vogl (2022) and Bianchi et al. (2023)). Readers are referred to Derman (2011) for an insightful exploration of the role of models in finance and to Ippoliti (2021) for some philosophical remarks on mathematics and finance.
The world of finance is always moving forward even in times of crisis. Innovations in finance come from the development of new financial services, products or technologies. Research trends in quantitative finance are driven not only by innovations, but also by structural changes in financial markets or by changes in regulation (Cesa (2017) and Carmona (2022)). When a structural change occurs some models are not anymore able to explain the phenomena observed in the market, consequently quants and researchers start working on new models. Examples of such changes are when the implied volatility smile appeared in 1987 (see Derman and Miller (2016)) or the Euribor-OIS spread materialized in 2007 (see Bianchetti and Carlicchi (2012)). Research activities driven by new products are, for instance, the development of pricing models for interest rate and equity derivatives started in ’90s, the structuring of credit products in the early 2000s, or the recent research trend on cryptocurrencies. New technologies applied to finance are namely the increasing role of big data and the advent of machine learning techniques. Regulation have an impact on the development of new quantitative tools for measuring, managing and monitoring financial risks (e.g. the Basel Accords).
In this paper we explore the arXiv preprint server, the dominant open-access preprint repository for scholarly papers in the fields of physics, mathematics, computer science, quantitative biology, quantitative finance, statistics, electrical engineering and systems science, and economics. The papers in arXiv are not peer-reviewed but there are advantages in submitting to this repository, mainly to disseminate a paper without waiting for the peer review and publishing process, which can be slow (see Huisman and Smits (2017)). The arXiv collection provides a unique source of data to conduct various studies, including bibliometric, trend, and citation network analyses (see Clement et al. (2019)). It is a valuable resource for advancing scientific knowledge and conducting research on research, often referred to as meta-research. For example, Eger et al. (2019) and Viet and Kravets (2019) perform trend detection on computer science papers stored in arXiv, Lin et al. (2020) conduct a case study of computer science preprints submitted to arXiv from 2008 to 2017 to quantify how many preprints have eventually been printed in peer-reviewed venues, Tan et al. (2021) explore the images of around 1.5 million of papers held in the repository, Okamura (2022) investigates the citations of more than 1.5 million preprints on arXiv to study the evolution of collective attention on scientific knowledge, Bohara et al. (2023) train a state-of-art classification approach, and Fatima et al. (2023) design an algorithm to help researchers to perform systematic literature reviews.
We study all papers on quantitative finance, a small portion of the entire arXiv containing more than two millions of works at the time of writing. The choice is also motivated by our experience in this domain and by scientific curiosity.
The code is run on a standard desktop environment, without recurring to a big cluster. Scaling to a large number of papers may be not trivial. Dealing with a large amount of data requires significant computing resources, including processing power and memory, to manipulate and analyze the data efficiently. It is not simple, and maybe even impossible, to explore more than two million of papers with a standard desktop environment like ours.
The studies of papers on finance topics is not new in the literature. Burton et al. (2020) review the history of a well-known journal in this field and highlight its growth in terms of productivity and impact. The authors present a bibliometric analysis and identify key contributors, themes, and co-authorship patterns and suggest future research directions. Ali and Bashir (2022) conduct a systematic literature review and a bibliometric analysis on around 3,000 articles on asset pricing sourced from the top 50 finance and economics journals, spanning a 47-year period from 1973 to 2020. As observed by the authors, the exclusion of certain publications may potentially offer an alternative perspective on the landscape of existing asset pricing research. By using bibliometric and network analysis techniques, including the Bibliometrix Tool of Aria and Cuccurullo (2017), Sharma et al. (2023) investigate more than 4,000 papers on option pricing appeared from 1973 to 2019. They follow the procedure suggested by Donthu et al. (2021). Their study aims to pinpoint high-quality research publications, to discern trends in research, to evaluate the contributions of prominent researchers, to assess contributions from different geographic regions and institutions, and, ultimately, to examine the interconnectedness among these aspects. The works of Burton et al. (2020), Ali and Bashir (2022) and Sharma et al. (2023) are focused on asset pricing or on a specific journal, their corpus is obtained by searching in the Scopus database some specific keywords, and the bibliometric analysis relies on VOS viewer (see Van Eck and Waltman (2010)) and Gephi (see Bastian et al. (2009)).
Our study explores all papers on quantitative finance collected in arXiv till the end of 2022 (around 16,000) and it considers text mining techniques implemented in Python to extract information directly from the portable document format (pdf) files containing the full text of the papers, excluding images, without relying on ad-hoc software or proprietary databases. Westergaard et al. (2018) found that examining the full text of documents significantly improved text mining compared to studies that only explored information collected from abstracts.222 As a crosscheck, we conduct the analysis on both abstracts and full texts. The analysis using the full text data shows better results. Their finding highlights the importance of using complete textual content for more comprehensive and accurate text mining and analysis.
The main objectives of our work are twofold. First, we explore the topics of the quantitative finance papers collected in arXiv in order to describe the evolution of topics over time. After having evaluated the performance of various clustering algorithms, we investigate on which themes researchers have focused their attention in the period from 1997 to 2022. Second, we try to understand who are the most prominent authors and journals in this field. Both analyses are performed with data mining techniques and without actually reading the papers.
The remainder of the paper is organized as follows. First, we provide a brief description of the data analyzed in this work (Section 2). Then, in Section 3 the preprocessing phase is discussed by offering further insights on the papers analyzed in our work. In Section 4 we compare various clustering algorithms and, after having selected the best performer, we explore, by splitting our corpus in 30 clusters, the evolution of topics over time. Finally, in Section 5 we describe an entity extraction process to investigate authors and journals with the largest number of occurrences in the corpus considered in this work. Section 6 concludes.
2 Data description
In this section we provide a description of the papers analyzed in this work. As observed above, there are various domains in arXiv (i.e. physics, mathematics, computer science, quantitative biology, quantitative finance, statistics, electrical engineering and systems science, and economics) and each domain has is own categories. The categories within the quantitative finance domain are the following:
  • computational finance (q-fin.CP) includes Monte Carlo, PDE, lattice and other numerical methods with applications to financial modeling;
  • economics (q-fin.EC) is an alias for econ.GN and it analyses micro and macro economics, international economics, theory of the firm, labor economics, and other economic topics outside finance;
  • general finance (q-fin.GN) is focused on the development of general quantitative methodologies with applications in finance;
  • mathematical finance (q-fin.MF) examines mathematical and analytical methods of finance, including stochastic, probabilistic and functional analysis, algebraic, geometric and other methods;
  • portfolio management (q-fin.PM) deals with security selection and optimization, capital allocation, investment strategies and performance measurement;
  • pricing of securities (q-fin.PR) discusses valuation and hedging of financial securities, their derivatives, and structured products;
  • risk management (q-fin.RM) is about risk measurement and management of financial risks in trading, banking, insurance, corporate and other applications;
  • statistical finance (q-fin.ST) includes statistical, econometric and econophysics analyses with applications to financial markets and economic data;
  • trading and market microstruture (q-fin.TR) studies market microstructure, liquidity, exchange and auction design, automated trading, agent-based modeling and market-making.
These categories are assigned by the authors when they submit their papers. Even if it is possible to select multiple couples of domain-category belonging to more than one domain, we select as reference category only the first category within the quantitative finance domain. Figure 1 shows the numbers of papers on quantitative finace submitted to arXiv between 1997 and 2022. The increase in the last three years is mainly due to the q-fin.EC category.
Figure 1: Categories by year.
As observed in Section 1, the code is implemented in Python and it is run under Ubuntu 22.04 on a desktop with an AMD Ryzen 5 5600g processor with 32GB of RAM. As we will describe in the following, numerous packages are considered.
As far as the collection process is concerned, we retrieve data from arXiv by selecting all categories within quantitative finance (i.e. q-fin). We collect articles metadata and pdf files for all articles from 1997 to 2022 for a total of around 16,000 articles (18GB of data).
While the metadata are obtained through urllib.request and feedparser, the pdf files are downloaded by means of the arxiv package. The metadata can be collected by following the suggestions provided in the arXiv web-pages. They are a fundamental input of the analysis and include the link to the paper main web-page, from where it is possible to extract the paper identification code (id, e.g. 2005.06390). The metadata contain information like authors names, paper title, primary category, submission and last update dates, abstract, and publication data when available (e.g. digital object identifier, DOI). Subsequent updates of the papers can be stored in the repository and for this reason there is the version number at the end of the paper id.
Since a paper could be assigned to multiple categories, a web-scraping tool written in Python allows us to retrieve from the paper main web-page the list of all categories of each single paper. We select from this list the subset of categories within the quantitative finance domain. Thus we assign as reference category of a paper the first category appearing in this subset. Starting from this list, we are able to filter and analyze all papers in the nine categories within q-fin.
The pdftotext package allows us to extract the text from pdf files. Each paper becomes a single (long) string. As discussed in Section 3, the length of these strings vary accross papers (see Figure 5), also because some documents are not papers (e.g. there are both theses and books). As a first assessment of the corpus, for each document we estimate the readability of the papers through textstat. As shown in Figure 2 the Flesch reading ease score (see DuBay (2004)) is on average equal to 65.7 (plain English), the lower and upper quartile are 59.91 (fairly difficult to read, but not far from the plain English) and 71.95 (fairly easy to read), respectively, and 99 per cent of the papers are in the range from 40.28 (difficult to read) and 88.20 (easy to read). There is only one paper with a negative value, but this is caused by the text contained in the figures. All other papers are above 17.17, that is above the extremely difficult to read level.
Figure 2: The Flesch reading ease score is reported. The vertical line represents the median, the dashed line is the quantile of level 0.25 (0.75), and the dotted line is the quantile of level 0.005 (0.995).
3 Text preprocessing
This section describes the text processing steps. The preprocessing phase is performed with nltk:(1) we split the text in tokens; (2) we extract the numbers representing years in the text;333 We assume these numbers have 4 digits. We do not explore this data in the empirical analysis. (3) we identify all strings containing alphabet letters, and we refer to them as words even in the case they do not belong to the English vocabulary; this step allows also to remove some symbols which are not recognized as letters in text analysis. Then, (4) we remove all stopwords and all words with lenght less than 3 characters; we also check whether there are words with more than 25 characters (quite uncommon in English); (5) we conduct a lemmatization by means of a part-of-speech tagger considering nouns, verbs, adjectives and adverbs; (6) we check if the paper is written in English by means of the langdetect package and we discard all non English papers. Both the extraction phase and the preliminary text analysis is parallelized by means of the multiprocessing package. We refer to the output of this first preprocessing phase as lemmatized data.
Figure 3: Frequent words of the corpus and their percentage of appearance.
Figure 4: Bigrams and tri(four)grams word clouds based on frequency with parameters min count equal to 250 and threshold equal to 10.
Thus, we analyze the frequency of the words across the whole corpus and we remove all words appearing less than 25 times. We discard also some words frequently used in writing papers on quantitative finance and which do not help in understanding the topic of the paper. The list of these words includes, for example, “proof” and “theorem”, verbs commonly used in mathematical sentences (e.g.,“assume”,“satisfy”, and “define”), mathematical functions (e.g. “min”, and “log”) and adverbs. The complete list is available upon request. In Figure 3 the list of the top 100 most frequent words obtained after this cleaning phase and their percentage of appearance is shown. The word “model” is extremely frequent (one every 100 words). The word “http” is also quite common, indicating that the papers full texts contain numerous internet links.
After a first preprocessing phase, we conduct an $n$-gram analysis by considering the Phrases model of the gensim package. We ignore all words and bigrams with total collected count over the entire corpus lower than 250 and set the score threshold equal to 10. We find the bigrams and then, to find trigrams and fourgrams, we apply again the same model to the trasformed corpus including bigrams. This approach allows us to have a better ex-post understanding of the corpus which is full of $n$-grams (e.g. Monte Carlo simulation, Eisenberg and Noe, or bank balance sheet). The wordclouds of the main bigrams and trigrams (fourgrams) are depicted in Figure 4.
It should be noted that some topic modeling algorithms analized in Section 4 do not need text preprocessing. In those cases the input is just a single list containing the whole paper text. While we refer to this latter input as raw data, we define the output of the preprocessing as cleaned data.
Figure 5: Papers length, in terms of number of words, for raw, lemmatized and cleaned data. The x-axis values are in thousands. The scale of x-axis varies across the three datasets. The vertical line represents the median, the dashed line is the quantile of level 0.25 (0.75), and the dotted line is the quantile of level 0.01 (0.99).
In Figure 5 we show how the length of the papers varies over the three data preprocessing phases. Starting from the raw data, containing all words and symbols, after a preliminary cleaning step we obtain the lemmatized data and then, after the last cleaning steps, the cleaned data. The number of words varies from a median value of 8,824 words for the raw data to around 2,518 words for the cleaned ones.
4 Topics trend
Now we are in the position to perform a topics trend analysis. We employ a topic modeling approach to identify the subjects discussed in the documents examined in this study, and then we observe how these topics change over time. Topic modeling refers to a class of statistical methods used to determine which subjects are prevalent in a given corpus. In Section 4.1 we select the best performing model among some selected approaches presented in the related literature. We evaluate these approaches by assessing their ability to accurately match the nine q-fin categories that researchers assign to their work when submitting it to arXiv. Then in Section 4.2, after having split the papers into 30 clusters, each one representing a specific topic, we discuss the evolution of research trends over time.
4.1 Algorithms performance on full texts
Topic modeling algorithms are widely used in natural language processing and text mining to uncover latent thematic structures in a collection of documents. Different algorithms have been developed, each with its own strengths and limitations (see Sethia et al. (2022)). The choice of a topic modeling algorithm depends on specific factors, such as the desired level of topic granularity and computational constraints. Some algorithms may require substantial computational resources and large amounts of training data. Here, we compare different algorithms by looking at some standard performance measures.
Since is not simple to assess the performance of different topic modeling algorithms (see Rüdiger et al. (2022)), we start by comparing the clusters assigned by each algorithm on the entire corpus to the nine clusters defined by the q-fin categories described in Section 2, that is the categories that researchers assign to their work during the submission process to arXiv. By exploiting a Bayesian optimization strategy, Terragni et al. (2021) present a framework for training, analyzing, and comparing topic models where the competitor models are trained by searching for their optimal hyperparameter configuration, for a given metric and dataset. Here we consider a simpler approach in which we compare the models by looking at some metrics. Evaluating topic modeling algorithms typically involves the use of performance measures, and it is important to note that different algorithms can yield varying results across these metrics.
We consider the following models.
  • K-means. We perform a clustering analysis by considering the K-means algorithm implemented in scikit-learn. This algorithm groups data points into K clusters by minimizing the distance between data points and their cluster center. The document word matrix is created through the CountVectorizer function, that converts the corpus to a matrix of token counts. We ignore terms that have a document frequency strictly higher than 75%. T
  • LDA. By considering the same document word matrix analyzed with the K-means algorithm, we perform topic modeling with the latent Dirichlet allocation (LDA). LDA is a well-known unsupervised learning algorithm. As observed in the seminal work of Blei et al. (2003), the basic idea is that documents are represented as random mixtures over latent topics, where each topic is characterized by a distribution over words. We study two different implementations of LDA (i.e. scikit-learn and gensim).
  • Word2Vec. We train a word embedding model (i.e. Word2Vec) and then we perform a clustering analysis by considering again the K-means approach. An embedding is a low-dimensional space into which high-dimensional vectors are projected. Machine learning on large inputs like sparse vectors representing words is easiser if embeddings are considered. Ideally, an embedding captures some of the semantics of the input by placing semantically similar inputs close together in the embedding space. The Word2Vec neural network introduces distributed word representations that capture syntactic and semantic word relationships (see Mikolov et al. (2013)). More in details, we generate document vectors using the trained Word2Vec model, that is, we get numerical vectors for each word in a document, and then the document vector is the weighted average of the vectors. Thus, the K-means algorithm is applied to the matrix representing the corpus. We consider the Word2Vec model implemented in gensim.
  • Doc2Vec. We create a vectorised representation of each document through the Doc2Vec model and then we perform a clustering analysis by considering the K-means approach. The Doc2Vec extends Word2Vec and it can learn distributed representations of varying lengths of text, from sentences to documents (see Le and Mikolov (2014)). We consider the Doc2Vec model implemented in gensim.
  • Top2Vec. We study the Top2Vec model, an unsupervised learning algorithm that finds topic vectors in a space of jointly embedded document and word vectors (see Angelov (2020)). This algorithm directly detects topics by performing the following steps. First, embedding vectors for documents and words are generated. Second, a dimensionality reduction on the vectors is implemented. Third, the vectors are clustered and topics are assigned. This algorithm is implemented in an ad-hoc library named Top2Vec and it automatically provides information on the number of topics, topic size, and words representing the topics.
  • BERTopic. We study a BERTopic model, which is similar to Top2Vec in terms of algorithmic structure and uses BERT as an embedder. As described in the seminal work of Grootendorst (2022), from the clusters of documents, topic representations are extracted using a custom class-based variation of term frequency-inverse document frequency (TF-IDF). This is the main difference with respect to Top2Vec. The algorithm is implemented in an ad-hoc library named BERTopic. The main downside with working with large documents, as in our case, is that information will be ignored if the documents are too long. A limited number of tokens are treated and anything longer is cut off. Since we are dealing with large documents, to work around this issue, we first split each documents in chunks of 300 tokens, thus we fit the model on these chunks. BERTopic does not allow one to directly select the number of topics, for this reason on the first step we obtain a number of topics much larger than the desired one. Since we obtain for each chunk the corresponding topic, we have for each document a list of possibly different topics and the length of these lists varies across documents (i.e. the length of a single list depends on the length of the corresponding document). To cluster this list of lists of topics, we consider each integer representing a topic as a word. Thus we use the Word2Vec algorithm described above to find similarities between these list of topics. Each topic label, that is the number representing the topic, is treated as a string, and Word2Vec transforms it into a numerical vector. We then apply K-means clustering to group these lists based on their similarity in the vector space. The resulting clusters reveal relationships and patterns among these lists and allow us to select the number of clusters we need for our purposes.
In theory both BERTopic and Top2Vec should use raw data since these algorithms rely on an embedding approach, and keeping the original structure of the text is of paramount importance (see Egger and Yu (2022)). However, raw data extracted from quantitative finance papers have a considerable amounts of formulas, symbols and numbers that may affect the algorithms performance. For this reason, we consider both raw and cleaned data. Additionally, for these state-of-the-art algorithms the number of extracted topics tends be large, but the algorithms offer the possibility to reduce the number of topics and of outliers, that can be larger than expected. The parameters of a BERTopic model have to be carefully chosen to avoid memory issues. Alternativelly, it is possible to perform an online topic modeling, that is the model is trained incrementally from a mini-batch of instances. This result in a less resource demanding approach in terms of memory and CPU usage, but it generates less rich and less comprehensive outputs and for these reasons we do not consider this incremental approach here.
In Table 1 we report the following similarity measures between true and predicted cluster labels: (1) the rand score (RS) is defined as the ratio between the number of agreeing pairs and the number of pairs, and it ranges between 0 and 1, where 1 stands for perfect match; (2) the adjusted rand score (ARS), that is the rand score adjusted for chances, has a value close to 0 for random labeling, independently of the number of clusters and samples, and exactly 1 when the clusterings are identical (up to a permutation), however is bounded below by -0.5 for especially discordant clusterings; (3) the mutual info score (MI) is independent of the absolute values of the labels (i.e. a permutation of the cluster labels does not change the value of the score); (4) the normalized mutual information (NMI) is a normalization of the MI to scale the results between 0 (no mutual information) and 1 (perfect correlation); (5) cluster accuracy (CA) is based on the Hungarian algorithm to find the optimal matching between true and predicted cluster labels; (6) to compute the purity score (PS), each cluster is assigned to the class which is most frequent in the cluster, and the similarity measure is obtained by counting the number of correctly assigned papers and dividing by the number of observations. This latter score increases as the number of clusters increases and for this reason, it cannot be used as a trade off between the number of clusters and clustering quality, that is to find the optimal number of clusters.
The measures presented in Table 1 demonstrates that the Doc2Vec approach, when coupled with K-means clustering on cleaned data, outperforms other models. This is evident from the higher MI and PS measures it achieves compared to its competitors. Moreover, Doc2Vec exhibits practical advantages, as it is straightforward to implement and significantly reduces computing time when compared to more advanced techniques like BERTopic. Interpreting the results of the Doc2Vec approach is simple, as it allows for the easy identification of the most representative documents by retrieving the centroid vectors of each cluster. The Word2Vec approach, when coupled with K-means clustering on cleaned data, shows also a good performance.
RS
ARS
MI
NMI
CA
PS
K-means
0.570
0.029
0.232
0.136
0.271
0.297
LDA scikit-learn
0.823
0.194
0.608
0.284
0.376
0.460
LDA gensim
0.788
0.085
0.276
0.131
0.275
0.314
Word2Vec K-means 0.832
0.200
0.613
0.283
0.371
0.427
Doc2Vec K-means
0.831
0.220
0.699
0.325
0.388
0.490
Top2Vec raw
0.810
0.195
0.501
0.239
0.365
0.404
Top2Vec cleaned
0.811
0.206
0.530
0.387
0.253
0.416
BERTopic raw
0.826
0.238
0.608
0.289
0.436
0.458
BERTopic cleaned
0.821
0.239
0.574
0.276
0.398
0.429
Table 1: Algorithms performance. We report the rand score (RS), the adjusted rand score (ARS), the mutual info score (MI), the normalized mutual info score (NIM), the cluster accuracy (CA), and the purity score (PS).
It is worth noting that there are no significant differences in performance measures when applying either the Top2Vec or BERTopic methods to raw or cleaned data. This could be attributed to the fact that raw data contain mathematical formulas that do not contribute substantial additional information, even if, as shown in Figure 5, raw data have a larger number of words. This finding indicating an equivalence between raw or cleaned data could be also attributed to the relatively simple structure commonly found in quantitative finance papers. It is important to note that these findings may not generalize to papers or books with more intricate and complex text structures and without formulas.
LDA implemented in scikit-learn has a better performance than the LDA implementation in gensim. The plain K-means does not show satisfactory results, even if the algorithm can be implement without a great effort.
Finally, as an overall assessment, it is important to highlight that the performance metrics reported in Table 1 do not demonstrate particularly impressive results. This could partly stem from the wide-ranging nature of each q-fin category, encompassing numerous subtopics and arguments. Conversely, some papers can be classified under multiple categories, as it is not always obvious how to select a single definitive category for a given work.
4.2 Empirical study
Figure 6: Topics trend by year across the sample of around 16,000 papers in the q-fin categories.
Figure 7: We show the topics clusters projected on the 2-dimensional space through the standard t-distributed stochastic neighbor embedding (t-SNE) algorithm. The numbers represent the topics reported in Table 2.
As shown in Section 4.1, the best performing model is Doc2Vec with K-means clustering applied on cleaned data. This model is considered to have a better understanding of the topics discussed in the quantitative finance papers analized in this work. To obtain the desired number of topics we perform again a K-means clustering analysis. To extract the most representative documents, we retrieve the centroid vectors of each cluster. These centroids represent the average position of all document embeddings assigned to a particular cluster. For each cluster centroid, we find the nearest neighbors among the original Doc2Vec embeddings. These nearest neighbors are the documents that are closest to the centroid in the embedding space and can be considered as the main documents of that cluster. Thus we select for each cluster the 20 most representative documents and we find a label for the topic on the basis of the documents titles. Note that the underlying meanings of the topics are subject to human interpretation. However, also this phase is automated by asking the topic to ChatGPT (GPT-3.5) after having provided the list of 20 titles (see Ebinezer (2023)).
Given the size of our sample, a reduction down to 30 topics can be considered a good compromise for a topic analysis. The selected number of topics strikes a good balance between ensuring a sufficient quantity of documents for each topic and maintaining the desired level of granularity. This approach allows us to extract meaningful insights from the data while avoiding an excessive division of content that could affect our ability to identify overarching trends and patterns. As shown in Figure 6, most of the topics have an increasing trend (topic 28 seems to be the only exception). This is also motivated by the growth in the number of papers shown in Figure 1. For each topic, the list of 20 titles is the input for the question we ask to the ChatGPT chatbot. The topics label (i.e. the ChatGPT reply to the question) together with the title of the most representative paper are reported in Table 2.
It is interesting to observe that topics related to decentralized finance and blockchain technology (2) and stock price prediction with deep learning and news sentiment analysis (20) show a remarkable increase. This is also true for health, policy, and social impact studies, represented by the topic 16, and diverse perspectives in education, innovation, and economic development (0). Both topics are oriented towards economics. These topics trends depend also by the introduction in 2014 of the q-fin.EC category within the arXiv quantitative finance papers. Classical quantitative finance subjects like portfolio optimization techniques and strategies (6), stochastic volatility modeling and option pricing (7), game theory and strategic decision-making (19) high-order numerical methods for option pricing in finance (23) as well as, new themes appeared in the literature in the last years, like deep reinforcement learning in stock trading and portfolio management (4) and environmental and economic impacts of mobility technologies (25) have attracted the interest of researchers in the analized period. The representativeness of topic 28 is limited, mainly because the number of papers in this cluster is low, and one could merge it with another cluster (i.e. 8).
It is clear that some topics are more related to economics than to finance. This also depends on the presence of the q-fin.EC category. For articles in this category there is not always a flawless alignment with quantitative finance.
To visualize the clusters, in Figure 7 the documents vectors are projected in the 2-dimensional space through the standard t-distributed stochastic neighbor embedding (t-SNE) algorithm. The larger the distance between topics, the more distinct the papers in those topics are in the original high-dimensional space. Conversely, it is also possible to have a better view on how close are topics each other (e.g. 8 and 28). It appears that the most specialized or narrowly-focused topics tend to occupy peripheral positions, while themes that are more aligned with economics are positioned closer to the center.
5 Extracting authors and journals
In this section we extract the author surnames by means of the spacy package. More in details, starting from the raw data, we perform a named-entities recognition. Since this approach extract both first names and surnames, we remove all first names by checking if these names are included in a list of about 67,000 first names. It should be noted that the number of occurence of the surname of the author in a paper strongly depends on the citation style. Surnames are always reported in the references, but they do not necessary appear in the main text of a paper. Additionally, even if the author-date style is widely used (i.e. the citation in the text consists of the authors name and year of publication), the surname of the first author appears more frequently (e.g. Bianchi is more probable than Tassinari, even if Bianchi and Tassinari are coauthors of the same papers, together with other coauthors).
The algorithm is able to find the names and surnames occuring in the text. These are included in the PERSON entity types. The first 100 authors by number of occurences in the corpus are selected. In order to have additional information on these authors, we obtain topics, number of citations, h-index and i10-index from Google Scholar (see Table 3). It should be noted that not all authors are registered in Google Scholar, even if they made a significant contribution to the field (e.g. Markowitz) or there are authors with the same surname and belonging to the same research field (see also Figure 3 in Sharma et al. (2023)). This is the case for some researchers we find in our corpus (e.g. Zhou, Bayraktar and Chakrabarti). For these last authors is not simple to find a perfect match in Google Scholar even if their number of occurrences is generally high.444 For the reasons described above we exclude from Table 3 the following researcher: Zhou, Markowitz, Peng, Jacod, Merton, Guo, Lo, Follmer, Yor, Almgren, Embrechts, Bayraktar, Artzner, Weber, Jarrow, Feng, Samuelson, Tang, Chakrabarti, Glasserman, Tsallis, Leung, Sato, Zariphopoulou, Kramkov, Karoui, Cizeau, Cao and Christensen.
The algorithm is also able to find the most cited journals, included in the ORG entity types: Journal of Finance (4490 $*{}^{*}$, Mathematical Finance (3785), Journal of Financial Economics ($*{}^{*}$, Physica A ($*{}^{*}$, Quantitative Finance (3044), Econometrica ($*{}^{*}$, Journal of Econometrics (1971), American Economic Review (1878), Insurance Mathematics and Economics (1667), Review of Financial Studies ($*{}^{*}$, Journal of Banking and Finance ($*{}^{*}$, Physical Review (1538), Journal of Economic Dynamics (1289), Energy (1267), Operations Research (1242), The Quarterly Journal of Economics ($*{}^{*}$, Journal of the American Statistical Association (1204), Management Science (1160), European Journal of Operational Research (1066), Quantum (1043), IEEE Transactions (996), Journal of Political Economy (990), Journal of Economic Theory (977), Energy Economics (946), International Journal of Theoretical and Applied Finance (946), SIAM Journal on Financial Mathematics (888), Science (865), Expert Systems with Applications (845), Applied Mathematical Finance (743), Finance and Stochastics (736), Mathematics of Operations Research (652), PLoS (602), The Annals of Applied Probability (593), Stochastic Processes and their Applications (525), Energy Policy (520), Internationl Journal of Forecasting (520), The European Physical Journal (514), Journal of Empirical Finance (510), Journal of Risk (509). We consider only journals with more than 500 occurrences and we esclude publishing houses. The journals with the asterisk symbol are identified with more than one name. It should be noted that some well-known journals are slightly below 500 occurences. Furthermore, it is worth noting that papers with a strong mathematical focus tend to receive significantly fewer citations compared to papers that lean more towards economics or finance.
It is important to acknowledge that while the arXiv repository serves as a valuable resource for scholarly papers, it may not encompass the entirety of the quantitative finance research landscape. While the repository strives to be inclusive and comprehensive, there may be variations in the representation of scholars from different countries. Some scholars may have a relatively higher presence due to their active participation in submitting their research to arXiv. The platform content is reliant on authors voluntarily submitting their work, which introduces the possibility of selection bias. As a result, some authors and their contributions may not be represented. Therefore, our analysis and conclusions should be interpreted within the context of the available arXiv data, recognizing that there may be additional research and authors in the field of quantitative finance who have chosen alternative avenues for publishing their work. The same observation is true for the findings described in Section 4.2.
It is possible that influential scholars may not be as consistently represented or that, for various reasons, they have not regularly submitted their work to arXiv (see also Metelko and Maver (2023)). In the study of Sharma et al. (2023), focused on option pricing, some well-known authors are cited but they do not appear among the first 100 authors in our analysis. This discrepancy could be influenced by factors such as publication preferences, possible copyright issues, or institutional practices that may vary across different academic communities.
As a final remark, in our view, researchers in quantitative finance should consider submitting their work to arXiv due to the potential benefits it offers (see also Mishkin et al. (2020)). The delay between arXiv posting and journal publication, which can sometimes be more than a year, underscores the importance of submitting preprints to the repository. By doing so, researchers can help the community to understand research trends in their field more promptly, while also accelerating the dissemination of their own findings. This approach aligns with the findings of Wang et al. (2020), who show that rapid and open dissemination through preprints helps scholarly and scientific communication, enhancing the reception of findings in the field. In the meanwhile, as a possible alternative to gain a comprehensive understanding of the entire landscape, future studies may consider incorporating other reputable academic databases and journals to ensure a more holistic exploration of quantitative finance research and its authors. Network approaches would also help to identify cliques and highly connected groups of authors. Insights from network clusters could further improve the understanding of the textual data investigated in this work.
6 Conclusions
In this study, we explore the field of quantitative finance through an analysis of papers in the arXiv repository. Our objectives are twofold: first, we investigate the evolution of topics over time, second, we identify prominent authors and journals in this domain. By employing data mining techniques, we achieve these goals without reading the papers individually.
The preprocessing phase, when needed, ensures the suitability of the data for subsequent analyses. Topic modeling helps in gaining insights and understanding the main themes and trends within our large dataset. By applying topic modeling algorithms, we identify the best performer and examine the temporal evolution of quantitative finance topics. This analysis reveals the changing research trends and highlights the emergence and decline of various topics over time.
Furthermore, we conduct an entity extraction process to identify influential authors and journals in the field. Through quantifying authors and journals occurrences, we shed light on the researchers who have made notable contributions to quantitative finance.
Our study demonstrates the power of data mining techniques in uncovering insights from a large-scale preprint repository. Our work not only showcases the power of data mining but also highlights the continued growth and dynamism of quantitative finance as a discipline. The techniques explored in this work can assist researchers in exploring and identifying novel research topics, discovering connections between different research areas, and staying up-to-date with the latest developments in the field. Furthermore, our methodology may serve as a roadmap for future studies on broader datasets or in other scientific domains utilizing text mining techniques. Although scaling to a larger number of papers may pose challenges, our approach provides valuable insights.
Finally, we believe that quantitative finance researchers should consider sharing their work on arXiv to potentially accelerate the dissemination and impact of their findings and to enhance the community understanding of research trends.